OpenStudy (anonymous):

the ratio of a map is 1: 200 000 two places are 71km apart. how far apart are they on the map??

OpenStudy (callisto):

1/ 200 000 = x/(71x100x1000) x = 35 cm If you understand the previous question, you should know how to do this!